OpenStudy (anonymous):

Find the slope of the line connecting the points (4,5) and (6,-1)

OpenStudy (anonymous):

I got \[\frac{ 3 }{ 1 }\] because i reduce please correct me if am wrong

OpenStudy (thunderbird336):

slope is change in y divided by change in x, so 5 - (-1) / 4 - 6 = 6 / -2 or -3

OpenStudy (thunderbird336):

m = (y2 - y1) / (x2 - x1)

OpenStudy (thunderbird336):

just be careful with negative numbers, it's easy to drop the minus sign...

OpenStudy (anonymous):

yes i got tht so far

OpenStudy (anonymous):

for this i got \[\frac{ -6 }{ 2}\]

OpenStudy (thunderbird336):

should be -1 - 5 in the numerator...

OpenStudy (thunderbird336):

but your -6 / 2 is good

OpenStudy (anonymous):

so do i reduce tht or no ?

OpenStudy (thunderbird336):

must've just been a typo cuz you got the right simplification and then the right answer I suppose

OpenStudy (thunderbird336):

yeah, you can and should I suppose... slope can be an integer

OpenStudy (anonymous):

ohh haha so i got tht correct thanks

OpenStudy (anonymous):

yeah i didnt know a slope can be a interger